OpenStudy (anonymous):

Rewrite the radical as a rational exponent. the fourth root of 7 to the fifth power

OpenStudy (anonymous):

wait im terrible in math and im doing a credit recovery class online . i need soo much help its not funny , can anyone help me ?

OpenStudy (the_fizicx99):

\(\ \sf \Large \sqrt[4]{7^5} \) ?

OpenStudy (the_fizicx99):

Is that what you mean? Well remember that \(\ \sf \Large x^{\frac{m}{n}} = \sqrt[n]{x^m}\)

OpenStudy (the_fizicx99):

So \(\ \sf \Large \sqrt[4]{7^5} = 7^{\frac{5}{4}}\)

OpenStudy (anonymous):

thank you sooo much . i love you lol

OpenStudy (the_fizicx99):

:3 Thanks, I know I'm kool c;

OpenStudy (the_fizicx99):

Here, this'll help you C;

OpenStudy (anonymous):

A rectangle has a length of the cube root of 81 inches and a width of 3 to the 2 over 3 power inches. Find the area of the rectangle. how would i multiply these two , to come up with my answer .